# Perception-Control Coupled Visual Servoing for Textureless Objects Using Keypoint-Based EKF

### v1

@@ Abstract
Visual servoing is fundamental to robotic applications, enabling precise positioning and control. However, applying it to textureless objects remains a challenge due to the absence of reliable visual features. Moreover, adverse visual conditions, such as occlusions, often corrupt visual feedback, leading to reduced accuracy and instability in visual servoing. In this work, we build upon learning-based keypoint detection for textureless objects and propose a method that enhances robustness by tightly integrating perception and control in a closed loop. Specifically, we employ an Extended Kalman Filter (EKF) that integrates per-frame keypoint measurements to estimate 6D object pose, which drives pose-based visual servoing (PBVS) for control. The resulting camera motion, in turn, enhances the tracking of subsequent keypoints, effectively closing the perception–control loop. Additionally, unlike standard PBVS, we propose a probabilistic control law that computes both camera velocity and its associated uncertainty, enabling uncertainty-aware control for safe and reliable operation. We validate our approach on real-world robotic platforms using quantitative metrics and grasping experiments, demonstrating that our method outperforms traditional visual servoing techniques in both accuracy and practical application.
@@ IV-A State Representation and Propagation
$\mathbf{F}_{k}={\begin{bmatrix}\frac{\partial\exp{\left(-\omega_{k}^{\wedge}\Delta t\right)}\mathbf{t}_{k}}{\partial\mathbf{t}_{k}}&\mathbf{0}_{3\times 3}\\ \mathbf{0}_{3\times 3}&\frac{\partial\text{Log}{\left(\exp{(-\bm{\omega}_{k}^{\wedge}\Delta t)}\mathbf{C}_{k}\right)}^{\vee}}{\partial\bm{\phi}_{k}}\end{bmatrix}}$, (8)
@@
$\frac{\partial\text{Log}{\left(\exp{(-\bm{\omega}_{k}^{\wedge}\Delta t)}\mathbf{C}_{k}\right)}^{\vee}}{\partial\bm{\phi}_{k}}=\mathcal{J}_{r}^{-1}\left(\text{Log}{\left(\exp{(-\bm{\omega}_{k}^{\wedge}\Delta t)}\mathbf{C}_{k}\right)}^{\vee}\right).$ (10)
Here, $\text{Log}\left(\cdot\right)^{\vee}$ denotes the logarithmic mapping from Lie group $\mathbb{SO}(3)$ to its Lie algebra $\mathfrak{so}(3)$. The operator $\mathcal{J}_{r}(\cdot)$ is the right Jacobian. This Jacobian is derived using the right perturbation model together with the Baker-Campbell-Hausdorff (BCH) approximation. For detailed derivations, readers are referred to [2].
@@ VI-C Evaluation Metrics
  • Trajectory Length Ratio (LR): Defined as the fraction of the method’s trajectory length divided by the length of the geodesic PBVS trajectory [11]. As illustrated in Figure 6(f), The geodesic of PBVS is the shortest path in SE(3) connecting the initial and desired camera poses, representing the most efficient trajectory under ideal conditions.
@@
TABLE I: Real-world performance comparison of visual servoing approaches. SR = Success Rate, TE = Translation Error, RE = Rotation Error, and LR = Length Ratio. Bold values indicate the best performance. Errors are reported as mean ± standard deviation over successful trials. Note that, for our approach, the results were obtained without activating the uncertainty-aware velocity reduction, allowing a fair comparison of standard servoing performance.

### v2

@@ Abstract
Visual servoing is fundamental to robotic applications, enabling precise positioning and control. However, applying it to textureless objects remains a challenge due to the absence of reliable visual features. Moreover, adverse visual conditions, such as occlusions, often corrupt visual feedback, leading to reduced accuracy and instability in visual servoing. In this work, we build upon learning-based keypoint detection for textureless objects and propose a method that enhances robustness by tightly integrating perception and control in a closed loop. Specifically, we employ an Extended Kalman Filter (EKF) that integrates per-frame keypoint measurements to estimate 6D object pose, which drives pose-based visual servoing (PBVS) for control. The resulting camera motion, in turn, enhances the tracking of subsequent keypoints, effectively closing the perception-control loop. Additionally, unlike standard PBVS, we propose a probabilistic control law that computes both camera velocity and its associated uncertainty, enabling uncertainty-aware control for safe and reliable operation. We validate our approach on real-world robotic platforms using quantitative metrics and grasping experiments, demonstrating that our method outperforms traditional visual servoing techniques in both accuracy and practical application.
@@ IV-A State Representation and Propagation
$\mathbf{F}_{k}={\begin{bmatrix}\frac{\partial\exp{\left(-\omega_{k}^{\wedge}\Delta t\right)}\mathbf{t}_{k}}{\partial\mathbf{t}_{k}}&\mathbf{0}_{3\times 3}\\ \mathbf{0}_{3\times 3}&\frac{\partial\text{ln}{\left(\exp{(-\bm{\omega}_{k}^{\wedge}\Delta t)}\mathbf{C}_{k}\right)}^{\vee}}{\partial\bm{\phi}_{k}}\end{bmatrix}}$, (8)
@@
$\frac{\partial\text{ln}{\left(\exp{(-\bm{\omega}_{k}^{\wedge}\Delta t)}\mathbf{C}_{k}\right)}^{\vee}}{\partial\bm{\phi}_{k}}=\mathcal{J}_{r}^{-1}\left(\text{ln}{\left(\exp{(-\bm{\omega}_{k}^{\wedge}\Delta t)}\mathbf{C}_{k}\right)}^{\vee}\right).$ (10)
Here, $\text{ln}\left(\cdot\right)^{\vee}$ denotes the logarithmic mapping from Lie group $\mathbb{SO}(3)$ to its Lie algebra $\mathfrak{so}(3)$. The operator $\mathcal{J}_{r}(\cdot)$ is the right Jacobian. This Jacobian is derived using the right perturbation model together with the Baker-Campbell-Hausdorff (BCH) approximation. For detailed derivations, readers are referred to [2].
@@ VI-C Evaluation Metrics
  • Trajectory Length Ratio (LR): Defined as the fraction of the method’s trajectory length divided by the length of the geodesic PBVS trajectory [11]. As illustrated in Figure 6(f), The geodesic of PBVS is the shortest path in SE(3) connecting the initial and desired camera poses, representing the most efficient trajectory.
@@
TABLE I: Real-world performance comparison of visual servoing approaches. Bold values indicate the best performance. Errors are reported as mean ± standard deviation over successful trials. Note that, for our approach, the results were obtained without activating the uncertainty-aware velocity reduction, allowing a fair comparison of standard servoing performance.
@@ VII CONCLUSION
Acknowledgements
This work was developed with the assistance of OpenAI’s GPT-4, which was used to (i) generate and refine code for robotic system integration, and (ii) editing and grammar enhancement for the manuscript. All AI-generated content was carefully reviewed and validated by the authors.